I have 2 liters of paint that is 30% yellow and 70% red. How many liters of yellow paint should I add to get a mixture that is 45% yellow and 55% red?
Answer by lwsshak3(11628)

I have 2 liters of paint that is 30% yellow and 70% red. How many liters of yellow paint should I add to get a mixture that is 45% yellow and 55% red.
**
let x=liters of yellow paint to add
x+2=liters of final mixture
..
total liters of yellow paint in mixture=30%*2+x=.6+x
total liters of red paint in mixture=70%*2=1.4
..
(x+2)*45%=.6+x
(x+2)=(.6+x)/45%
..
(x+2)*55%=1.4
(x+2)=1.4/55%
..
(.6+x)/45%=1.4/55%
.45*1.4=.55*(.6+x)
.63=.33+.55x
.55x=.30
x=.30/.55=.5455
ans:
liters of yellow paint to add=.5455 liters